Two cars start in a race with velocities u1 and u2 and travel in a st.line with accelerations xand y. If both reach finish line at same time ,the range of the race is 
  1. 2(u1-u2)/(x-y)2 X(u1y-u2x)
  2. 2(u1-u2)/(x+y) X(u1x-u2y)
  3. 2(u1-u2)2/(x-y)2
  4. 2u1u2/xy

Grade:11

1 Answers

noogler
8 years ago
solve it by using s=ut+1/2at2
 
as both cars travel same dis. in same time
 s=u1t+1/2xt2=u2t+1/2yt2
solve them for t........................u get t=2(u1-u2)/(x-y)2 X(u1y-u2x)
 
APPROVE IF USEFUL!!!!!!!!!!